# VF-Plan: Bridging the Art Gallery Problem and Static LiDAR Scanning with Visibility Field Optimization

### v1

@@ 4.2 Visibility Field
Each viewpoint is assigned a View Angle, with the scene represented as a grid where each cell center is a viewpoint. Thus, the scene is represented as a Visibility Field, as shown in Figure 3(b), which aligns well with LiDAR properties and expert experience, displaying expected visibility patterns.
@@ 4.3 Converging Lines and Joints
Observation data naturally converge to the medial axis and then joints of the Visibility Field, so placing viewpoints along these skeletons and joints ensures full coverage. Thus, VPN optimization reduces from a 2D to a 1D problem along these skeletons. We use an medial axis algorithm [43] to identify skeletons and joint points, placing candidate viewpoints along these lines. Where overlap ratios are insufficient, additional viewpoints are added to ensure network cohesion.
For any skeleton ridge $R_{ij}$ connecting joints $J_{i}$ and $J_{j}$, we calculate an Overlap Ratio $O_{ij}$ of the endpoints. When $O_{ij}$ falls below a threshold, a midpoint $J_{m}$ is added to split the ridge, ensuring connectivity. Both the skeleton joints and connecting points are called as converging points. And the resulting skeleton ridges are called converging lines. This iterative placement process continues until all converging points have adequate overlap for registrability.
@@ 4.4 Overlap Ratio
Figure 5: Overlap calculation between observed segments from different viewpoints, quantifying connectivity.
The Overlap Ratio quantifies shared visibility between viewpoints, a requirement for robust registration. Five overlap ratios based on length and angle are defined, ensuring effective connectivity independent of distance alone.
Given wall segments observed by viewpoint $V_{a}$ as $L_{a}$ with an observation angle $\theta_{a}$:
@@
The length-based overlap ratios are:
@@
And the angle-based overlap ratios are:
@@ 4.5 Greedy Optimization for VPN Construction
The VF-Plan greedy algorithm constructs an efficient Viewpoint Network (VPN) by starting with candidate viewpoints positioned strategically at key joints within the Visibility Field. For each pair of candidate viewpoints, an Overlap Ratio $O$ is calculated, with pairs showing adequate overlap retained as adjacencies in the VPN to ensure robust connectivity.
@@
To systematically track visibility, we construct the Completeness Table $C$ (see Equation 1), which determines whether a given viewpoint $v_{i}$ observes a specific line segment $l_{j}$. To enhance precision in observation tracking, wall boundaries are uniformly partitioned into equal-length segments.
@@
    Initialize Completeness Table $C$
@@
As outlined in Algorithm 1 and illustrated in Figure 6, the greedy optimization initiates by selecting the viewpoint with the highest coverage as the seed. The algorithm then iteratively expands by adding neighboring viewpoints that maximize additional coverage, ensuring each new viewpoint significantly contributes to covering line segments while constructing a minimal, fully connected Viewpoint Network (VPN).
@@ 5.1 Dataset
In this section, we evaluate the proposed VPN optimization algorithm across multiple scenarios and compare its performance with state-of-the-art methods in both indoor and outdoor environments.
Due to the lack of a dedicated benchmark for VPN optimization, we compiled a dataset from established sources, focusing on two primary scene types: indoor and outdoor environments. This dataset provides a robust foundation for evaluating VPN optimization across a range of realistic settings.
For indoor environments, we selected scenes with varied room layouts to evaluate the algorithm’s adaptability across multiple configurations. The Structure3D dataset [56] was included for its extensive collection of diverse Asian residential scenes, featuring both complex and compact room types frequently used in 3D modeling and synthetic data generation. VPNs generated from this dataset can also facilitate synthetic LiDAR data production for advancing research in 3D scene understanding. Specific scenes such as 00007, 00009, 00071, 02970, and 02972, discussed in subsequent sections, are drawn from Structure3D. For further comparisons, we incorporated additional scenes from Zeng1 [51], Noichl [34], Xu [48], and TUB1 from the ISPRS benchmark [52, 24].
For outdoor scenarios, we limited our approach to 2D site plans without considering terrain elevation, distinguishing our method from elevation-based approaches like [8]. Selected scenes include the CCIT and Crowsnest datasets from the University of Calgary [20] and the ECUT dataset from East China University of Technology [8], which represent open-space environments suitable for VPN optimization.
@@ 5.2 Performance Metrics
The optimization algorithm aims to construct a minimal, fully connected viewpoint network with reduced redundancy. As most state-of-the-art methods achieve 100% coverage, we focus on two key metrics: the number of optimal viewpoints and network connectivity.
Network connectivity is measured using the Weighted Average Path Length (WAPL), which reflects the compactness of the network [5]. A shorter WAPL indicates stronger connectivity and a more efficient structure. For a network with weighted edges $w$, WAPL is calculated as:
@@
5.3 Algorithm Parameter Effects
5.3.1 Overlap Ratio
We propose five overlap ratios according to the length of scanned walls and according angles to viewpoints in section 4.4. Here we evaluate them in two test indoor scenes, illustrated in Figure 7. Among them, $O^{\text{Min\_Len}}$ quickly reaches 100%, showing limited variability along movement paths, as seen from B to H in Figures 7(b) and (e), making it less suitable for dynamic scenes. The angle-based measures, $O^{\text{Union\_Ang}}$ and $O^{\text{Mean\_Ang}}$, exhibit non-monotonic behavior, particularly along paths from C to E, shown in Figures 7(c) and (f), which limits their usefulness for consistent registration. The $O^{\text{Union\_Len}}$ measure tends to be too small, often underestimating connectivity. Therefore, we select $O^{\text{Mean\_Len}}$ for subsequent experiments and analyses due to its balanced performance across different scenarios.
@@ 5.3.2 Scan Radius
For outdoor scenarios, we examine the impact of varying $R_{\text{max}}$, testing values at 15m, 30m, 45m, and 75m, with $R_{\text{min}}$ fixed at 1m. As illustrated in Figure 8, increasing $R_{\text{max}}$ expands the area of valid placements for viewpoints—defined where the View Angle exceeds 1.0—and reduces the Viewpoint Count (VC) needed for optimized coverage, from 23 to 14, 10, and 8. Larger $R_{\text{max}}$ values are advantageous for open spaces, while smaller values require additional viewpoints for full coverage. These results indicate that our method aligns effectively with real-world scanner specifications and expert guidelines.
Figure 8: Effect of $R_{\text{max}}$ on Viewpoint Placement and Network Efficiency. As $R_{\text{max}}$ increases from 15m to 75m, the valid placement area (where View Angle $>$ 1.0) grows, reducing the Viewpoint Count (VC) needed for optimized VPN coverage from 23 to 8.
5.3.3 Wall Partition
To improve station coverage accuracy, walls are partitioned into segments of varying lengths, as shown in Figure 9. We find that partitioning does not alter the final optimization result, but only affects the selection order of stations. Larger partition lengths (or no partitioning) place initial viewpoints at major intersections, optimizing computational cost and maintaining effective coverage.
@@ 5.3.4 Effect of Windows
Including windows in the model adds complexity to the skeleton structure, generating more detailed skeletons with additional joint points, as shown in Figure 10. When windows are considered, the VPN algorithm requires only a minimal increase in viewpoints to achieve complete coverage, as seen in Table 1.
Without windows, the Viewpoint Count (VC) averages around 1.4 times the number of rooms, while with windows, it increases slightly to 1.6 times the room count. These results indicate that our algorithm effectively accommodates the added detail from windows. For precise window scans, including window frames in the optimization process is recommended.
Figure 10: Comparison of VPN structures with and without windows. Including windows adds complexity to the skeleton and slightly increases the Viewpoint Count.
@@
Table 1: Viewpoint Count (VC) comparison for room coverage with and without windows across different scenes.
@@ 5.4 Results and Analysis
Zeng1 0.47 0.53 43.37 12 12 10 Zeng[51]
Noichl 0.32 0.38 23.91 8 10 7 Noichl2023[34]
TUB1 1.03 1.00 40.97 28 29 19 Zhai2024[52]
Xu 0.35 0.31 28.69 8 8 7 Xu2017[48]
CCIT 0.52 0.47 38.12 8 8 7 Jia2018[20]
Crowsnest 0.50 0.49 34.02 9 8 8 Jia2018[20]
ECUT 0.67 0.76 11.94 58 52 139 Chen2023[8]
Table 2: Comparison of viewpoint counts (VC) and Weighted Average Path Length (WAPL) across indoor and outdoor scenes. Lower WAPL indicates more compact network connectivity.
We benchmarked our method against state-of-the-art approaches using datasets from prior studies, representing a range of indoor and outdoor scenes. Indoor datasets include Zeng1 [51], Noichl [34], and TUB1 from the ISPRS dataset [52]. Outdoor datasets consist of CCIT and Crowsnest from the University of Calgary [20] and ECUT from East China University of Technology [8].
Indoor Results. As shown in Figure 12, our method’s viewpoint optimization closely aligns with expert-designed networks, achieving comprehensive coverage and creating compact, interconnected structures. This optimized network design enhances registration accuracy and connectivity, even in complex indoor settings.
Outdoor Results. For outdoor scenes, our method effectively adapts to open spaces, as illustrated in Figure 13. Adjustments to the $R_{\text{max}}$ parameter notably improved performance in expansive environments, reducing the required number of viewpoints without compromising coverage.
Table 2 demonstrates that our method performs comparably or better than expert-designed networks, achieving similar viewpoint counts (VC) and lower Weighted Average Path Length (WAPL). This outcome indicates a more compact and cohesive network structure. Unlike most baselines that often yield fragmented networks (highlighted by green ellipses in Figures 12 and 13), our approach maintains low WAPL scores, which reflect enhanced connectivity and minimized redundancy.
In indoor environments, our method achieves a 98% reduction in WAPL compared to baseline methods, creating highly connected networks well-suited for confined spaces. For complex outdoor scenarios, such as ECUT, our approach achieves a 91% WAPL reduction over baselines, demonstrating scalability and adaptability across both intricate indoor layouts and expansive outdoor scenes.
To further validate the optimized viewpoints, we simulated static LiDAR scanning at each optimized viewpoint using Helios++ [13]. As shown in Figure 11, the simulated LiDAR point clouds confirm complete scene coverage, with effective overlap between adjacent scans across both indoor and outdoor environments.
@@ 6 Conclusions
We introduced a novel solution to the Viewpoint Planning Problem and Art Gallery Problem in static LiDAR scanning by leveraging a Visibility Field to reduce optimization complexity from 2D to 1D. This VF-based approach enables the creation of a minimal, fully connected viewpoint network that ensures comprehensive coverage with reduced redundancy, advancing viewpoint planning for static LiDAR applications.

### v2

@@ 4.2 Visibility Field (VF)
Each viewpoint is assigned a View Angle, with the scene represented as a grid where each cell center is a viewpoint. Thus, the scene is represented as a VF, as shown in fig.˜3(b), which aligns well with LiDAR properties and expert experience, displaying expected visibility patterns.
Figure 3: Visibility Field (VF) for static LiDAR scanning. (a, c) VF computed with only one wall (green highlight); (b, d) VF with all walls, revealing the global visibility structure. (a, b) show a rectangular room, and (c, d) depict a flat layout.
@@ 4.3 Converging Lines and Joints
In the VF, observation data naturally converge toward the medial axis and its joint points. Placing viewpoints along these skeletal structures guarantees full coverage (section˜3, appendix˜A), reducing VPN optimization from a 2D search to a 1D problem. We extract the medial axis using [SP08] via OpenCV [Bra00] and detect joints by restricting inside room polygons: a skeleton point is marked as a joint candidate if at least three of its eight neighbors are also skeleton points. Connected-component analysis groups candidates, and each cluster’s centroid is retained as a representative joint.
For each skeleton ridge $R_{ij}$ between joints $J_{i}$ and $J_{j}$, we compute the overlap ratio $O_{ij}$. If $O_{ij}$ is below the threshold, a midpoint $J_{m}$ is inserted to split the ridge, ensuring connectivity. We refer to all joints and added points as converging points, and their connecting ridges as converging lines. This process iterates until every converging point satisfies the overlap requirement for reliable registration.
@@ 4.4 Overlap Ratio
Figure 4: Overlap computation between wall segments from two viewpoints. (a, b) Segments visible from viewpoints A and B. (c) Segments jointly visible, used to measure connectivity.
The overlap ratio quantifies shared visibility between two viewpoints and is key for robust registration. We define five variants based on either segment length or angular alignment, evaluated for their ability to maintain connectivity independent of viewpoint spacing.
As shown in fig.˜4, $L_{a}$ and $L_{b}$ denote wall segments visible from viewpoints $V_{a}$ and $V_{b}$, with observation angles $\theta_{a}$ and $\theta_{b}$:
@@
Length-based ratios:
@@
Angle-based ratios, where $\theta_{a}^{ab}$ and $\theta_{b}^{ab}$ are the angular spans of $L_{ab}$:
@@ 4.5 Greedy Optimization for VPN Construction
The VF-Plan greedy algorithm constructs an efficient VPN by initializing candidate viewpoints at strategically selected joint points within the VF. For each pair of candidate viewpoints, an Overlap Ratio $O$ is computed to quantify shared visibility. Pairs exceeding a predefined overlap threshold are retained as edges in the VPN, ensuring robust connectivity. To enforce the coverage constraint, we construct a binary Coverage Table $C$ (see eq.˜1) that records whether each viewpoint observes each structural segment.
@@
    Initialize Coverage Table $C$
@@
As outlined in Algorithm 1 and illustrated in fig.˜5, the greedy optimization initiates by selecting the viewpoint with the highest coverage as the seed. The algorithm then iteratively expands by adding neighboring viewpoints that maximize additional coverage, ensuring each new viewpoint significantly contributes to covering line segments while constructing a minimal, fully connected VPN.
@@ 5 Experiments
In this section, we evaluate the proposed VPN optimization algorithm across multiple scenarios and compare its performance with state-of-the-art methods in both indoor and outdoor environments. Due to the lack of a dedicated benchmark for VPN optimization, we compiled a dataset from established sources, focusing on two primary scene types: indoor and outdoor environments. This dataset provides a robust foundation for evaluating VPN optimization across a range of realistic settings. The pipeline is implemented in C++14 and compiled on Windows 11, running on an Intel Core i9 CPU (24 cores, 2.20 GHz) with 32 GB RAM.
@@ 5.1 Dataset
For indoor environments, we selected scenes with varied room layouts to evaluate the algorithm’s adaptability across multiple configurations. The Structure3D dataset [ZZL20] was included for its extensive collection of diverse Asian residential scenes, featuring both complex and compact room types frequently used in 3D modeling and synthetic data generation. VPNs generated from this dataset can also facilitate synthetic LiDAR data production for advancing research in 3D scene understanding. Specific scenes such as 00007, 00009, 00071, 02970, and 02972, discussed in subsequent sections, are drawn from Structure3D. For further comparisons, we incorporated additional scenes from Zeng1 [ZLC22], Noichl [NB23], Xu [XZY17], and TUB1 from the ISPRS benchmark [ZZG24, KTA21].
For outdoor scenarios, we limited our approach to 2D site plans without considering terrain elevation, distinguishing our method from elevation-based approaches like [CZH22]. Selected scenes include the CCIT and Crowsnest datasets from the University of Calgary [JL18] and the ECUT dataset from East China University of Technology [CZH22], which represent open-space environments suitable for VPN optimization.
@@ 5.2 Performance Metrics
The optimization algorithm aims to construct a minimal, fully connected VPN with reduced redundancy. As most state-of-the-art methods achieve 100% coverage, we focus on two key metrics: the number of optimal viewpoints and network connectivity.
Network connectivity is measured using the Weighted Average Path Length (WAPL), which reflects the compactness of the network [BLM06]. A shorter WAPL indicates stronger connectivity and a more efficient structure. For a network with weighted edges $w$, WAPL is calculated as:
@@
5.3 Ablation Study
We evaluate overlap ratio type, scan radius, window inclusion, and, in appendix˜B, wall partition, VF grid resolution and overlap thresholds. VPN performance is largely insensitive to wall partition length and VF grid resolution, though these parameters affect computation time, while others require tuning for specific scenarios. Default settings for subsequent experiments are: $O^{\text{Mean\_Len}}$ for overlap computation, overlap ratio $0.4$ indoors and $0.3$ outdoors, $R_{\min}=0.6$ m indoors and $1.2$ m outdoors, $R_{\max}=30.0$ m indoors and $75.0$ m outdoors, wall partition $0.1$ m indoors and $1.0$ m outdoors, and VF grid resolution $0.02$ m indoors and $0.25$ m outdoors.
5.3.1 Overlap Ratio Type
We propose five overlap ratios according to the length of scanned walls and according angles to viewpoints in section˜4.4. Here we evaluate them in two test indoor scenes, illustrated in fig.˜6. Among them, $O^{\text{Min\_Len}}$ quickly reaches 100%, showing limited variability along movement paths, as seen from B to H in fig.˜6(b) and (e), making it less suitable for dynamic scenes. The angle-based measures, $O^{\text{Union\_Ang}}$ and $O^{\text{Mean\_Ang}}$, exhibit non-monotonic behavior, particularly along paths from C to E, shown in fig.˜6(c) and (f), which limits their usefulness for consistent registration. The $O^{\text{Union\_Len}}$ measure tends to be too small, often underestimating connectivity. Therefore, we select $O^{\text{Mean\_Len}}$ for subsequent experiments and analyses due to its balanced performance across different scenarios.
Figure 7: Effect of $R_{\text{max}}$ on Viewpoint Placement and Network Efficiency. As $R_{\text{max}}$ increases from 15m to 75m, the valid placement area (where View Angle >> 1.0) grows, reducing the VC needed for optimized VPN coverage from 23 to 8.
@@ 5.3.2 Scan Radius
For outdoor scenarios, we examine the impact of varying $R_{\text{max}}$, testing values at 15m, 30m, 45m, and 75m, with $R_{\text{min}}$ fixed at 1m. As illustrated in fig.˜7, increasing $R_{\text{max}}$ expands the area of valid placements for viewpoints—defined where the View Angle exceeds 1.0—and reduces the Viewpoint Count (VC) needed for optimized coverage, from 23 to 14, 10, and 8. Larger $R_{\text{max}}$ values are advantageous for open spaces, while smaller values require additional viewpoints for full coverage. These results indicate that our method aligns effectively with real-world scanner specifications and expert guidelines.
Figure 8: Comparison of VPN structures with and without windows. Including windows (bottom row) adds complexity to the skeleton and slightly increases the VC.
@@ 5.3.3 Effect of Windows
Incorporating windows into the model increases the complexity of the skeleton structure by introducing additional joint points, resulting in a more detailed medial axis, as illustrated in fig.˜8. However, accurate scanning of windows is often essential in practical applications. To assess the impact of including windows, we evaluate the performance of our algorithm with and without windows considered in the optimization. When windows are included, the VPN algorithm requires only a slight increase in the number of viewpoints to achieve complete coverage, as shown in table˜1. On average, the number of viewpoints is approximately 1.4 times the number of rooms without windows, and increases modestly to 1.6 times with windows. These results demonstrate that the algorithm effectively handles the additional geometric detail introduced by windows. For applications requiring precise window coverage, including window frames in the optimization process is recommended.
@@
Table 1: VC comparison for room coverage with and without windows across different scenes.
@@ 5.4 Results and Analysis
Zeng1 0.47 0.53 43.37 12 12 10 Zeng[ZLC22]
Noichl 0.32 0.38 23.91 8 10 7 Noichl2023[NB23]
TUB1 1.03 1.00 40.97 28 29 19 Zhai2024[ZZG24]
Xu 0.35 0.31 28.69 8 8 7 Xu2017[XZY17]
CCIT 0.52 0.47 38.12 8 8 7 Jia2018[JL18]
Crowsnest 0.50 0.49 34.02 9 8 8 Jia2018[JL18]
ECUT 0.67 0.76 11.94 58 52 139 Chen2023[CZH22]
Table 2: Comparison of VC and WAPL across indoor and outdoor scenes. Lower WAPL indicates more compact network connectivity.
We benchmarked our method against state-of-the-art approaches using datasets from prior studies, representing a range of indoor and outdoor scenes. Indoor datasets include Zeng1 [ZLC22], Noichl [NB23], and TUB1 from the ISPRS dataset [ZZG24]. Outdoor datasets consist of CCIT and Crowsnest from the University of Calgary [JL18] and ECUT from East China University of Technology [CZH22].
Indoor Results. As shown in fig.˜9, our method’s viewpoint optimization closely aligns with expert-designed networks, achieving comprehensive coverage and creating compact, interconnected structures. This optimized network design enhances registration accuracy and connectivity, even in complex indoor settings.
Outdoor Results. For outdoor scenes, our method effectively adapts to open spaces, as illustrated in fig.˜10. Adjustments to the $R_{\text{max}}$ parameter notably improved performance in expansive environments, reducing the required number of viewpoints without compromising coverage.
Table˜2 shows that our method matches or surpasses expert-designed networks while significantly outperforming baseline methods. Across all scenes, our VC remains comparable to experts, differing by at most one viewpoint, yet our WAPL is consistently lower in five of seven cases, indicating more compact and cohesive networks. Compared to baselines, the improvement is substantial: in indoor scenes such as Zeng1 and Noichl, WAPL is reduced by 98% and 97% respectively, eliminating the fragmented connectivity seen in baseline outputs (green ellipses in figs.˜9 and 10). In challenging outdoor cases, reductions remain high, including 91% in ECUT and 87% in CCIT, demonstrating scalability to large and complex environments. These results confirm that our VF-based optimization preserves expert-level coverage while achieving markedly superior network compactness over existing automated methods.
To further validate the optimized viewpoints, we simulated static LiDAR scanning at each optimized viewpoint using Helios++ [EYW22]. As shown in fig.˜11, the simulated LiDAR point clouds confirm complete scene coverage, with effective overlap between adjacent scans across both indoor and outdoor environments.